OpenStudy (anonymous):

Prove that \(\cos 72=\frac{\phi}{2}\)

OpenStudy (anonymous):

where \(\phi\) is the \[\color{brown}{Golden \space Ratio}\]

ganeshie8 (ganeshie8):

\[\large x = e^{i \dfrac{360}{5}t } , ~~0\leq t\leq 4\] are the five roots of polynomial \(\large x^5-1\)

ganeshie8 (ganeshie8):

we can work out the roots by factoring \(\large x^5-1 \) and comparing it with the five roots of unity complete work is here http://www.integraltec.com/math/math.php?f=cosine72.html

OpenStudy (anonymous):

YES YES ACTUALLY THE REAL ORIGINAL PROBLEM IS FINDING THE ROOTS OF UNITY \[z_k=\cos \frac{k\theta }{n}+i \sin \frac{k\theta}{n} \] \[0\le n\le k-1\] for k=5

OpenStudy (anonymous):

i was wondering if i can compute cos 72 without a calculator,somehow

ganeshie8 (ganeshie8):

Ahh I see :) the algebraic method shown in that site is pretty neat - they are dividing x-1 and find finding the remainign four roots by factoring

OpenStudy (anonymous):

very NEAT solution,thank you @ganeshie8 apperiated

ganeshie8 (ganeshie8):

here is another way to work it using simple trig http://mathforum.org/library/drmath/view/53992.html this is also looking good to me if we just want the value of cos(72) :)

ganeshie8 (ganeshie8):